OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 5. Find the number(s).

OpenStudy (anonymous):

Write this as an equation, where the number = x

OpenStudy (anonymous):

im not good at writing equations @IvoryKnight

OpenStudy (anonymous):

You are going to need to learn if you are going to get anywhere in math. At the higher levels, everything is an equation. Ok lets look at this one number at a time, on one side, you are adding a number and 5 times that number's reciprocal. On the other side, you have 5. Try writing out the equation, if you make a mistake, I will at least know what direction to point you in.

OpenStudy (anonymous):

As a general tip, "sum", always means "put in a plus sign" and "is" always means "="

OpenStudy (anonymous):

6x+5=x

OpenStudy (shayaan_mustafa):

Hi tbrooks3 :) How are you dear? :)

OpenStudy (shayaan_mustafa):

Let us take the required number as "x"

OpenStudy (anonymous):

hi im fine, thanks for asking. how r u?

OpenStudy (shayaan_mustafa):

The sum of a number and 6 times its reciprocal is 5. The sum of a number and 6 times its reciprocal = 5 The sum of "x" and 6 times "1/x" = 5 The x + 6(1/x) = 5 Finally, \[\Large x+\frac{6}{x}=5\]

OpenStudy (shayaan_mustafa):

I am fine too :)

OpenStudy (shayaan_mustafa):

Have you got how did I get the equation?

OpenStudy (anonymous):

yes i solved it and recieved the 2 solutions (2,3)... thanks alot :)

OpenStudy (shayaan_mustafa):

Yes great.. Ever welcome.

OpenStudy (anonymous):

@Shayaan_Mustafa if i want to find A number minus 20 times its reciprocal is 1. would the equation be \[x -\left(\begin{matrix}20 \\ x\end{matrix}\right)=1\]

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

thanks

OpenStudy (shayaan_mustafa):

Yes. You got it. Now I can say my labor worked. lol

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

im stuck again . would i just do \[50\div7\] for the sum of an integer and its reciprocal is 50/7. Find the integer.

OpenStudy (shayaan_mustafa):

Let say that required integer is "x" The sum of an integer and its reciprocal is 50/7 The sum of an integer and its reciprocal = 50/7 The sum of "x" and "1/x" = 50/7 x + 1/x = 50/7 Finally, \[\Large x+\frac{1}{x}=\frac{50}{7}\]

OpenStudy (anonymous):

ohh okay thanks

OpenStudy (shayaan_mustafa):

My pleasure :D

OpenStudy (anonymous):

could the solution be \[\left(\begin{matrix}1 \\ 7\end{matrix}\right) , 7\]

OpenStudy (anonymous):

or would i simplify the fraction?

OpenStudy (shayaan_mustafa):

You can check your answer by substituting each in the given equation. And if your answer is right then right hand side and left hand side of the equation will be same.